OpenStudy (anonymous):

Solve the following equation y^2=13/36

OpenStudy (anonymous):

so you solving for y, right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

\[y^2- \frac{ 13 }{ 36 }= 0 \rightarrow y^2-\frac{ 13 }{ 36 } = 0\]

OpenStudy (anonymous):